# Spectral Bias Mitigation via xLSTM-PINN: Memory-Gated Representation Refinement for Physics-Informed Learning

### v1

@@ 2.2 Spectral-Bias Suppression and Frequency-Domain Gain: Kernel-Tail Lifting in xLSTM-PINN
Eq. (26) states the exponential advantage in endpoint error; Eq. (27) shows the reduction of time-to-threshold; Eq. (28) presents the condition for advancing the resolvable bandwidth. We supply three details in Sec IV of Supplementary Materials (SM): (i) linearization error bounds for Eqs.(17)–(20) and a sufficient smallness condition on $\|A\|$; (ii) a verifiable necessary-and-sufficient criterion for the monotonicity of $\alpha(\mathbf{k})$, based on Rayleigh-quotient ordering of $B$ along feature directions $v_{\mathbf{k}}$; (iii) an upper bound for the gap between $\exp(SA)$ and $(I+A)^{S}$, which tightens Eq.(24) via $\sigma_{\min}^{2}\!\bigl((I+A)^{S}\bigr)$.
Fig. 2 matches these claims one-to-one: panel Fig. 2 (a) reports a marked drop of $E_{T}(|\mathbf{k}|)=\|u_{T}-\phi_{\mathbf{k}}\|_{L^{2}}$ at mid-to-high frequencies; panel Fig. 2 (b) shows the broadband gain $G(|\mathbf{k}|)=E_{T}^{\mathrm{base}}(|\mathbf{k}|)/E_{T}^{\mathrm{xLSTM}}(|\mathbf{k}|)$ with confidence bands; panel Fig. 2 (c) displays the downward shift of $\tau(|\mathbf{k}|)$ and the rightward motion of the front $k^{*}(\varepsilon)$. These frequency-domain benchmarks verify that gated memory and residual micro-steps upgrade the single-layer effective map to $(I+A)^{S}$, lift high-frequency kernel weights, mitigate spectral bias, and improve accuracy under the same training budget.
@@ 3 Numerical Experiments
3.1 1D Advection–Reaction (Drift–Decay)
We study the constant-coefficient first-order advection–reaction equation
$u_{t}+a\,u_{x}+b\,u=0,\qquad[x,t]\in[0,2]\times[0,1],$ (29)
and choose $a=-\tfrac{1}{2}$, $b=\tfrac{1}{2}$. The inflow boundary lies on the right edge; we impose
$u[x,0]=6\,\mathrm{e}^{-3x},\qquad u[2,t]=6\,\mathrm{e}^{-6-2t}.$ (30)
The analytic solution
$u^{*}[x,t]=6\,\mathrm{e}^{-3x-2t}$ (31)
satisfies the PDE and both constraints exactly. We use physics-constrained learning to approximate $u^{*}$. Let $u_{\theta}[x,t]$ denote the network output and define the PDE residual
We keep the sampling sizes consistent with this section: $|\mathcal{S}_{r}|=3000$ and $|\mathcal{S}_{t=0}|+|\mathcal{S}_{x=2}|=500$. On the representation side, we replace only the stack with xLSTM blocks that combine gated memory and residual micro-steps; according to the frequency-domain gain analysis in Sec 2, this design increases the effective convergence weights along high-frequency directions and thus mitigates spectral bias.
Fig. 3 reports the results. The first row (xLSTM-PINN) shows near-perfect overlap between the diagonal decay bands of the reference and the prediction; the absolute-error map remains deep blue over the whole domain, with only a thin, mild stripe aligned with the characteristic direction. The second row (baseline PINN) exhibits large warm-colored regions in the interior, a significantly thickened error band along the characteristic direction, and block-like hot spots near the bottom edge and the upper-right corner. We therefore conclude that xLSTM-PINN suppresses high-wavenumber components along characteristics more rapidly, yielding a colder and more uniform error field, consistent with the spectral-bias mitigation mechanism in Sec 2.
Model MSE RMSE MAE MaxAE
xLSTM–PINN 6.28e-06 2.51e-03 1.54e-03 1.71e-02
PINN 5.62e-05 7.50e-03 7.44e-03 9.69e-03
Tab. 1 quantifies the comparison: xLSTM-PINN reduces MSE, RMSE, and MAE substantially relative to the baseline, while MaxAE is slightly higher due to a localized corner peak and does not dominate the global error surface. We present the training loss histories in Fig. 4; they further confirm the improved convergence behavior of xLSTM-PINN.
@@ 3.3 Non-dimensional Steady-State Heat Conduction in a Circular Silicon Plate with Convective Boundary
We start from the dimensional heat–transfer model. We assume the silicon wafer is thin enough to justify a two–dimensional approximation, and we define the physical temperature field $T(X,Y)$ on the disk
@@
Using the ambient temperature $T_{\infty}$ as the reference, we set the temperature rise $\Delta T=T-T_{\infty}$. With constant conductivity $k$ and a uniform volumetric heat source $Q\,[\mathrm{W/m}^{3}]$, we write the governing equation and the Newton cooling boundary:
@@
We nondimensionalize the model. We choose the characteristic length $R$ and the temperature scale $T_{\mathrm{ref}}=\dfrac{QR^{2}}{k}$, and define
$x=\frac{X}{R},\qquad y=\frac{Y}{R},\qquad\theta(x,y)=\frac{\Delta T(X,Y)}{T_{\mathrm{ref}}},\qquad\Omega=\{(x,y):\,x^{2}+y^{2}\leq 1\}.$
Using $\partial_{X}=(1/R)\partial_{x}$ and $\partial_{Y}=(1/R)\partial_{y}$, we obtain the dimensionless model
@@
where the Biot number $\mathrm{Bi}=hR/k$. We adopt the same physical parameters and boundary settings as the finite–element benchmark in [tao2025lnn], and we use its solution of FEM as a pointwise validation set.
Let $\theta_{\theta}(x,y)$ denote the network approximation. We define the interior residual
$r_{\theta}(x,y)=\theta_{\theta,xx}(x,y)+\theta_{\theta,yy}(x,y)+1,$
and the boundary residual
$b_{\theta}(x,y)=-\nabla\theta_{\theta}(x,y)\!\cdot n(x,y)-\mathrm{Bi}\,\theta_{\theta}(x,y).$
We minimize a weighted empirical $\ell_{2}$ objective over interior samples $\mathcal{S}_{\Omega}$ and boundary samples $\mathcal{S}_{\partial\Omega}$:
@@
Automatic differentiation supplies first– and second–order derivatives. We match the implementation’s sampling sizes: $|\mathcal{S}_{\Omega}|=3000$ in the interior and $|\mathcal{S}_{\partial\Omega}|=500$ on the boundary.
In Fig. 7 we compare predictions with the FE validation set. xLSTM-PINN produces concentric contour lines that align with the analytic parabolic profile; the absolute–error map stays cold across the disk and forms only an ultrathin ring near the rim, with magnitude on the order of $10^{-4}$. The baseline PINN develops a wider and brighter boundary ring that diffuses inward into slowly undulating curved errors, at magnitudes one to two orders larger. We attribute this gap to the frequency–domain mechanism established in Sec 2: gated memory and residual micro–steps raise effective convergence weights along high–frequency directions, markedly suppress spectral bias, and therefore satisfy the Poisson source and the Robin flux more accurately under the same training budget.
In Tab. 3 we report four metrics consistent with this error structure: xLSTM-PINN reduces MSE from $2.86\times 10^{-7}$ to $9.66\times 10^{-9}$, RMSE from $5.35\times 10^{-4}$ to $9.83\times 10^{-5}$, MAE from $5.26\times 10^{-4}$ to $7.87\times 10^{-5}$, and MaxAE from $7.46\times 10^{-4}$ to $4.03\times 10^{-4}$. The training loss histories, presented in Fig. 8, further validate the improved convergence characteristics of xLSTM-PINN.
@@ 3.4 Anisotropic Poisson–Beam Equation
We solve on the unit square $\Omega=(0,1)^{2}$ the fourth–order mixed operator
$u_{xx}-u_{yyyy}=f(x,y),\qquad f(x,y)=(2-x^{2})\,\mathrm{e}^{-y},$ (42)
and impose the boundary conditions
$u(x,0)=x^{2},\quad u_{yy}(x,0)=x^{2},\quad u(x,1)=\frac{x^{2}}{\mathrm{e}},\quad u_{yy}(x,1)=\frac{x^{2}}{\mathrm{e}},$ (43)
$u(0,y)=0,\quad u(1,y)=\mathrm{e}^{-y}.$ (44)
We validate with the analytic solution
$u^{*}(x,y)=x^{2}\,\mathrm{e}^{-y}.$ (45)
We specify sampling sets and their cardinalities. We draw the interior residual set $S_{r}\subset\Omega$ by i.i.d. uniform sampling with $|S_{r}|=N=1000$. On the boundary, we sample
@@
and set the same size for each set, $|S_{\bullet}|=N_{\partial}=1000$. In this way, we enforce the PDE residual in the interior, match $u$ and $u_{yy}$ on the two horizontal edges, and impose Dirichlet data on the two vertical edges.
Using automatic differentiation, we obtain $u_{\theta,xx}$, $u_{\theta,yyyy}$, and $u_{\theta,yy}$ for the network output $u_{\theta}$. We minimize the weighted squared objective
@@
We generate a validation set from $u^{*}$ and compute MSE, RMSE, MAE, and MaxAE on a dense grid; Tab. 4 summarizes these metrics. The visual comparison in Fig. 9 shows that xLSTM-PINN produces a cleaner upper-right corner and transition bands near boundaries, achieves markedly lower peak amplitudes than the baseline PINN, and attenuates high-frequency error ripples. These observations confirm that we effectively suppress spectral bias and improve overall accuracy for this stiff, high-order constrained problem. We keep $N$ and all boundary set sizes identical to the values stated above.

### v2

@@ 2.2 Spectral-Bias Suppression and Frequency-Domain Gain: Kernel-Tail Lifting in xLSTM-PINN
establishes the expansion of learnable frequency range under precision tolerance $\varepsilon$.
Supplementary Materials (SM) Section IV provides complete technical elaboration of these results: (i) linearization error bounds for Eqs. (17)–(20) with sufficient conditions on $\|A\|$; (ii) verifiable necessary-and-sufficient criterion for $\alpha(\mathbf{k})$ monotonicity via Rayleigh-quotient ordering; (iii) refined bound tightening Eq. (24) through $\sigma_{\min}^{2}((I+A)^{S})$ analysis.
Experimental measurements in Fig. 2 provide comprehensive validation of these theoretical predictions. Panel (a) confirms systematic error reduction at mid-high frequencies, panel (b) demonstrates broadband spectral gain exceeding unity with statistical significance, and panel (c) exhibits the characteristic downward shift in convergence time with rightward bandwidth extension. These frequency-domain measurements conclusively verify that gated memory and residual micro-steps transform the layer mapping to $(I+A)^{S}$, elevate high-frequency kernel eigenvalues, suppress spectral bias, and enhance accuracy under identical training constraints.
We evaluate xLSTM-PINN performance on the two-dimensional Laplace equation with mixed boundary conditions, a canonical benchmark in potential theory that rigorously tests numerical methods’ ability to handle diverse constraint types. The governing equation for the potential field $\phi(x,y)$ is specified as:
$\Delta\phi=\phi_{xx}+\phi_{yy}=0,\quad(x,y)\in[0,1]\times[0,1],$ (29)
subject to mixed Dirichlet–Neumann boundary conditions:
$\phi(x,0)=0,\quad\phi(x,1)=1,\quad\partial_{x}\phi(0,y)=0,\quad\partial_{x}\phi(1,y)=0.$ (30)
This boundary value problem admits the analytic solution $\phi^{*}(x,y)=y$, providing an exact reference for quantitative accuracy assessment.
The neural network approximation $\phi_{\theta}(x,y)$ is trained through minimization of a composite objective function that simultaneously enforces the governing physics and boundary constraints. Defining the interior residual as:
$r_{\theta}(x,y)=\phi_{\theta,xx}(x,y)+\phi_{\theta,yy}(x,y),$ (31)
we construct the complete optimization objective:
$\begin{split}\mathcal{J}(\theta)=&\lambda_{\Omega}\frac{1}{|\mathcal{S}_{\Omega}|}\sum_{(x,y)\in\mathcal{S}_{\Omega}}\left|r_{\theta}(x,y)\right|^{2}+\lambda_{D0}\frac{1}{|\mathcal{S}_{y=0}|}\sum_{x\in\mathcal{S}_{y=0}}\left|\phi_{\theta}(x,0)\right|^{2}\\ &+\lambda_{D1}\frac{1}{|\mathcal{S}_{y=1}|}\sum_{x\in\mathcal{S}_{y=1}}\left|\phi_{\theta}(x,1)-1\right|^{2}+\lambda_{N0}\frac{1}{|\mathcal{S}_{x=0}|}\sum_{y\in\mathcal{S}_{x=0}}\left|\partial_{x}\phi_{\theta}(0,y)\right|^{2}\\ &+\lambda_{N1}\frac{1}{|\mathcal{S}_{x=1}|}\sum_{y\in\mathcal{S}_{x=1}}\left|\partial_{x}\phi_{\theta}(1,y)\right|^{2}.\end{split}$ (32)
Automatic differentiation enables exact computation of the required first and second-order derivatives for residual evaluation. To ensure fair comparison, we maintain consistent sampling densities across all experiments, utilizing 1000 interior collocation points with an additional 1000 points distributed along each boundary segment.
Visual analysis in Fig. 3 reveals profound qualitative differences in solution fidelity between the proposed and baseline methodologies. The xLSTM-PINN prediction demonstrates exceptional agreement with analytic isocontours, with absolute errors consistently maintained at approximately $3.8\times 10^{-4}$ throughout the computational domain. These minimal discrepancies manifest as fine vertical striations that approach the limits of numerical precision. Conversely, the baseline PINN generates a smooth, curved error band of order $10^{-2}$ that permeates the entire domain, exhibiting the characteristic low-frequency shape bias inherent in conventional neural architectures.
Quantitative metrics presented in Table 1 provide rigorous numerical validation of these observational findings. The xLSTM-PINN achieves substantial error reduction across all evaluation criteria: MSE diminishes from $6.04\times 10^{-5}$ to $1.47\times 10^{-8}$, RMSE from $7.77\times 10^{-3}$ to $1.21\times 10^{-4}$, MAE from $7.75\times 10^{-3}$ to $9.90\times 10^{-5}$, and MaxAE from $8.81\times 10^{-3}$ to $3.82\times 10^{-4}$. These comprehensive improvements directly correlate with the spatial error distributions observed in Fig. 3 and substantiate the theoretical mechanism established in Section 2, where gated memory and residual micro-steps enhance high-frequency convergence to drive numerical errors toward machine precision levels.
Complementary examination of training dynamics in Fig. 4 offers additional verification, demonstrating that xLSTM-PINN achieves accelerated convergence to significantly lower loss plateaus compared to the baseline implementation. This enhanced optimization behavior further exemplifies the practical advantages of spectral-bias suppression through architectural innovation, confirming that xLSTM-PINN delivers superior performance while preserving the fundamental principles of physics-constrained learning.
@@ 3.2 Non-dimensional Steady-State Heat Conduction in a Circular Silicon Plate with Convective Boundary
We examine the performance of xLSTM-PINN on a steady-state heat conduction problem characterized by internal heat generation and convective boundary conditions, representing a physically relevant scenario in thermal engineering applications. The analysis begins with the dimensional heat transfer formulation, considering a thin silicon wafer that justifies two-dimensional approximation. The physical temperature field $T(X,Y)$ is defined on the circular domain:
@@
Referencing the ambient temperature $T_{\infty}$, we define the temperature rise $\Delta T=T-T_{\infty}$. With constant thermal conductivity $k$ and uniform volumetric heat source $Q\,[\mathrm{W/m}^{3}]$, the governing equation and Newton cooling boundary condition are expressed as:
@@
To establish a dimensionless formulation, we select characteristic length $R$ and temperature scale $T_{\mathrm{ref}}=QR^{2}/k$, defining the normalized variables:
$x=\frac{X}{R},\quad y=\frac{Y}{R},\quad\theta(x,y)=\frac{\Delta T(X,Y)}{T_{\mathrm{ref}}},\quad\Omega=\{(x,y):x^{2}+y^{2}\leq 1\}.$ (39)
Applying the coordinate transformations $\partial_{X}=(1/R)\partial_{x}$ and $\partial_{Y}=(1/R)\partial_{y}$ yields the dimensionless boundary value problem:
@@
where the Biot number $\mathrm{Bi}=hR/k$ characterizes the relative importance of convective to conductive heat transfer. We maintain identical physical parameters and boundary configurations to the finite-element benchmark established in [tao2025lnn], utilizing the FEM solution as a pointwise validation reference.
The neural network approximation $\theta_{\theta}(x,y)$ is optimized through minimization of a composite objective function that enforces both the interior physics and boundary conditions. The interior residual captures satisfaction of the Poisson equation:
$r_{\theta}(x,y)=\theta_{\theta,xx}(x,y)+\theta_{\theta,yy}(x,y)+1,$ (41)
while the boundary residual incorporates the convective flux condition:
$b_{\theta}(x,y)=-\nabla\theta_{\theta}(x,y)\cdot n(x,y)-\mathrm{Bi}\,\theta_{\theta}(x,y).$ (42)
The complete optimization objective combines these components:
@@
Automatic differentiation provides exact computation of the required first and second-order spatial derivatives. Consistent with the implementation specifications, we employ 3000 interior sampling points and 500 boundary points to ensure comprehensive constraint enforcement.
Comparative analysis in Fig. 5 demonstrates substantial differences in solution quality between the proposed and baseline methods. The xLSTM-PINN generates concentric isothermal contours that closely align with the reference parabolic profile, with absolute errors confined to an ultrathin boundary ring of magnitude approximately $10^{-4}$. In contrast, the baseline PINN produces a significantly wider and more intense error ring that penetrates inward with slowly undulating patterns, exhibiting error magnitudes one to two orders larger. This discrepancy directly reflects the spectral bias mitigation capabilities established in Section 2, where gated memory and residual micro-steps enhance high-frequency convergence to achieve superior boundary condition satisfaction under identical computational resources.
Quantitative metrics presented in Table 2 provide rigorous numerical confirmation of these observations. The xLSTM-PINN achieves comprehensive error reduction across all evaluation criteria: MSE decreases from $2.86\times 10^{-7}$ to $9.66\times 10^{-9}$, RMSE from $5.35\times 10^{-4}$ to $9.83\times 10^{-5}$, MAE from $5.26\times 10^{-4}$ to $7.87\times 10^{-5}$, and MaxAE from $7.46\times 10^{-4}$ to $4.03\times 10^{-4}$. Complementary analysis of training dynamics in Fig. 6 further validates these findings, demonstrating that xLSTM-PINN achieves faster convergence to lower loss plateaus, thereby confirming the practical optimization benefits arising from architectural spectral bias suppression.
@@ 3.3 Anisotropic Poisson–Beam Equation
We investigate the performance of xLSTM-PINN on the anisotropic Poisson-Beam equation, a challenging fourth-order partial differential equation that tests numerical methods’ capacity to handle high-order derivatives and mixed boundary conditions. The governing equation defined in the unit square domain $\Omega=(0,1)^{2}$ is expressed as follows:
$u_{xx}-u_{yyyy}=f(x,y),\qquad f(x,y)=(2-x^{2})e^{-y},$ (44)
subject to comprehensive boundary constraints that include both function values and second derivatives:
$u(x,0)$ $=x^{2},$ $u_{yy}(x,0)$ $=x^{2},$ (45)
$u(x,1)$ $=\frac{x^{2}}{e},$ $u_{yy}(x,1)$ $=\frac{x^{2}}{e},$
$u(0,y)$ $=0,$ $u(1,y)$ $=e^{-y}.$
This boundary value problem admits the analytic solution:
$u^{*}(x,y)=x^{2}e^{-y},$ (46)
providing an exact reference for rigorous accuracy assessment.
The computational implementation employs structured sampling strategies to ensure comprehensive constraint enforcement throughout the domain. The interior residual set $S_{r}\subset\Omega$ contains $N=1000$ points generated through independent uniform sampling. Boundary conditions are enforced through dedicated sampling sets:
@@
with each boundary set containing $N_{\partial}=1000$ points. This sampling strategy ensures balanced enforcement of the governing PDE, second-derivative matching on horizontal boundaries, and Dirichlet conditions on vertical boundaries.
The network approximation $u_{\theta}(x,y)$ is optimized through minimization of a comprehensive objective function that incorporates all physical constraints. Automatic differentiation provides exact computation of the required derivatives $u_{\theta,xx}$, $u_{\theta,yyyy}$, and $u_{\theta,yy}$. The complete optimization objective is formulated as:
@@
Visual analysis in Fig. 7 reveals significant qualitative differences in solution accuracy between the proposed and baseline methods. The xLSTM-PINN generates solutions with clean transition regions near boundaries, particularly in the upper-right corner, while maintaining markedly lower error amplitudes throughout the domain. The absolute error distribution exhibits weak ripples with substantially attenuated high-frequency oscillations. In contrast, the baseline PINN produces pronounced error patterns characterized by broad undulations and distinct stripe artifacts, indicating unresolved high-frequency components and spectral bias limitations.
Quantitative evaluation against the analytic reference solution confirms these observational findings. Comprehensive error metrics presented in Table 3 demonstrate consistent improvement across all measured criteria, with xLSTM-PINN achieving superior accuracy in MSE, RMSE, MAE, and MaxAE. These results substantiate the effective suppression of spectral bias even for this stiff, high-order constrained problem, where conventional architectures struggle to resolve the complex solution structure. The maintained sampling consistency with $N=1000$ interior points and equivalent boundary set sizes ensures fair comparison and highlights the architectural advantages of xLSTM-PINN in handling challenging differential operators.
@@
3.4 1D Advection–Reaction (Drift–Decay)
We examine the performance of xLSTM-PINN on the first-order advection-reaction equation, a fundamental model for transport phenomena with decay processes. The governing equation is defined as:
$u_{t}+a\,u_{x}+b\,u=0,\qquad[x,t]\in[0,2]\times[0,1],$ (49)
with constant coefficients $a=-\tfrac{1}{2}$ and $b=\tfrac{1}{2}$ representing advection velocity and reaction rate respectively. The inflow boundary condition is specified on the right edge due to the negative advection velocity, with complete boundary constraints:
$u[x,0]=6\,\mathrm{e}^{-3x},\qquad u[2,t]=6\,\mathrm{e}^{-6-2t}.$ (50)
This boundary value problem admits the analytic solution:
$u^{*}[x,t]=6\,\mathrm{e}^{-3x-2t}$ (51)
which exactly satisfies both the governing equation and boundary conditions, providing a rigorous benchmark for accuracy assessment.
The neural network approximation $u_{\theta}(x,t)$ is trained through physics-constrained learning, with the residual defined as:
Consistent sampling densities are maintained with $|\mathcal{S}_{r}|=3000$ interior points and a total of 500 boundary points distributed between initial and boundary condition enforcement. The architectural modification is confined to the representation level, where standard MLP layers are replaced with xLSTM blocks that integrate gated memory and residual micro-steps, thereby enhancing high-frequency convergence according to the spectral analysis established in Section 2.
Visual analysis in Fig. 9 reveals substantial differences in solution quality and error distribution between the compared methods. The xLSTM-PINN prediction exhibits near-perfect alignment with the characteristic decay pattern of the reference solution, maintaining minimal absolute error throughout the domain. The error distribution manifests as a thin, mild stripe aligned with the characteristic direction, consistent with the expected propagation pattern. In contrast, the baseline PINN displays significant error accumulation characterized by a thickened band along characteristics, widespread warm-colored regions in the interior domain, and localized hot spots near the bottom edge and upper-right corner, indicating unresolved high-frequency components.
Model MSE RMSE MAE MaxAE
xLSTM–PINN 6.28e-06 2.51e-03 1.54e-03 1.71e-02
PINN 5.62e-05 7.50e-03 7.44e-03 9.69e-03
Quantitative evaluation presented in Table 4 substantiates these visual observations with rigorous numerical evidence. The xLSTM-PINN achieves significant error reduction across primary metrics: MSE decreases by an order of magnitude from $5.62\times 10^{-5}$ to $6.28\times 10^{-6}$, RMSE from $7.50\times 10^{-3}$ to $2.51\times 10^{-3}$, and MAE from $7.44\times 10^{-3}$ to $1.54\times 10^{-3}$. The marginally higher MaxAE in xLSTM-PINN arises from a localized corner peak that does not dominate the global error distribution, reflecting the method’s capacity to suppress widespread error propagation while maintaining overall accuracy.
Complementary analysis of training dynamics in Fig. 10 provides additional validation, demonstrating that xLSTM-PINN achieves accelerated convergence to lower loss plateaus compared to the baseline. This enhanced optimization behavior, coupled with the observed error reduction patterns, confirms the practical benefits of spectral-bias suppression through architectural innovation, establishing xLSTM-PINN as an effective approach for resolving challenging transport phenomena.